Head-First Company plans to sell 5,000 baseball bats at $75 each in the coming year. Unit contribution margin is $25. Total fixed cost equals $49,500 (includes fixed factory overhead and fixed selling and administrative expense). Operating income at 5,000 units sold is $75,500.

The degree of operating leverage (rounded to the nearest tenth) is
Answer #1

Total Contribution margin = 5,000*25 = 125,000

Net operating income = Total Contribution margin - Fixed cost

= 125,000 - 49,500 = 75,500

Degree of operating leverage = Contribution margin/Net operating income

= 125,000/75,500

= 1.66